OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The perimeter of a certain rectangle is no more than 94 feet. If the length is eleven feet more than the width, what is the width?

OpenStudy (cwtan):

inequality?

OpenStudy (anonymous):

I have no idea! The question was very confusing to me

OpenStudy (anonymous):

5-Step Problem Solving Method Familiarize Translate Carry out Check State

OpenStudy (cwtan):

"Non 5-step method" == x is length y is width 2(x+y)<94 x=11+y x-11=y 2(x+(x-11))<94 4x-22<94 4x<116 x<29